2) Cola Co. manufactures a product with a standard direct labor cost of two hours at $24.00 per hour. During July, 2,000 units were produced using 4,200 hours at $24.40 per hour. The labor efficiency variance was?

A. $4,880 Favorable

B. $4,800 Unfavorable

C. $4,800 Favorable

Thanks! :-)
Answer #1

Labor efficiency variance = (SH-AH) * SR

= (2000*2 - 4200)*24

= 200*24

= 4800 UNFAVORABLE

Option B :)